OpenStudy (anonymous):

write the expression as a single natural logarithm 2 in x-4in c

OpenStudy (unklerhaukus):

do you mean \[2\ln x-4\ln c\]?

OpenStudy (unklerhaukus):

\[n\log(x)=\log(x^n)\] \[\log(y)+\log(z)=\log(yz)\]

OpenStudy (unklerhaukus):

get those coefficient into the indices, then use the product property

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

ok so the answer would be... in (x^2-c^4)

OpenStudy (unklerhaukus):

*Whops its a difference of logs, use the division property \[\log(x)-\log(y)=\log(x/y)\]

OpenStudy (anonymous):

answer---> in x^2/c^4 or in C^4/x^2

OpenStudy (anonymous):

idk which one lol

OpenStudy (unklerhaukus):

\[2\ln x-4\ln c=\ln x^2-\ln c^4=\ln (\cdots/\cdots)\]

OpenStudy (anonymous):

in(x^2/C^4)

OpenStudy (unklerhaukus):

right, in fact the division property is a special case of the product property \[\log(x)-\log(y)=\log(x)+\log(y^{-1})=\log(x)+\log(1/y)=\log(x/y)\]

OpenStudy (anonymous):

thank you it makes more sense